OpenStudy (anonymous):

If I have an equation: 7X/2 +5=8 How would I get X? Do I divide the fraction or multiply?

OpenStudy (anonymous):

7X/2=3

OpenStudy (shaik0124):

7x/2=8-5 7x/2=3 7x=6 x=6/7

OpenStudy (anonymous):

X=6/7

OpenStudy (anonymous):

Oh so I would multiply the 2 and 3, then put the 6 and 7 into a fraction?

OpenStudy (anonymous):

How would that be a fraction, I'm lost on that part

OpenStudy (anonymous):

what part

OpenStudy (anonymous):

The part about the fraction at the end?

OpenStudy (anonymous):

\[7x/2 +5=8\] \[7x/2=3\] \[7x=6\] \[x=6/7\]

OpenStudy (anonymous):

6/7?

OpenStudy (anonymous):

thats your answer for x

OpenStudy (anonymous):

its about 0.86

OpenStudy (shaik0124):

i solved already no need to rply same thing@ld

OpenStudy (anonymous):

well i did to @shaik0124

OpenStudy (shaik0124):

y

OpenStudy (anonymous):

cause i can

OpenStudy (shaik0124):

:)

OpenStudy (anonymous):

Thank You both for the help :)

OpenStudy (shaik0124):

now go to sleep or take rest :)